what is the solution of this equation 7x+3=3x-17? thank you! god bless you !
Answer by rothauserc(4718)

7x +3 = 3x -17
subtract 3x from both sides of =
4x +3 = -17
subtract 3 from both sides of =
4x = -20
divide both sides of = by 4
x = -5
check the answer by using original equation
7*-5 + 3 = 3*-5 -17
-35 +3 = -15 -17
-32 = -32
our answer checks :-)